# Superadditive Communication with the Green Machine: A Practical Demonstration of Nonlocality without Entanglement .1 Abstract

### v3

@@
Superadditive Communication with the Green Machine: A Practical Demonstration of Nonlocality without Entanglement
@@
Achieving the ultimate Holevo limit of optical communication capacity requires a joint-detection receiver which makes a collective quantum measurement over multiple modulated symbols. Such superadditivity—a higher communication rate than that achievable by symbol-by-symbol optical detection—is a special case of the well-known nonlocality without entanglement and has yet to be demonstrated. In this article, we propose and demonstrate a design of joint-detection receivers, the Green Machine, that can achieve superadditivity. We build this receiver and show that its capacity surpasses any symbol-by-symbol receivers in the photon-starved regime with binary-phase-shift-keying (BPSK). Our Green Machine receiver can also significantly reduce the transmitter peak power requirement compared with the pulse-position modulation (the conventional modulation format used for deep space laser communication). We further show that the self-referenced phase makes it immune to phase noise, e.g., atmospheric turbulence or platform vibrations.
.2 Introduction
@@
The fundamental limit of reliable optical communication rate, i.e., capacity, is governed by the laws of quantum mechanics and formulated as the Holevo-Schumacher-Westmoreland (HSW) theorem—colloquially known as the Holevo capacity [1, 2]. Holevo capacity is higher than the Shannon capacity, which is achievable by symbol-by-symbol detection. This gap becomes more prominent at the photon-starved regime where the received photon flux is low [3, 4]. To close the gap, a joint-detection receiver applying a collective measurement across a codeword comprising multiple symbols is necessary. For a given encoding format, the fact that joint detection can achieve a bits-per-symbol rate that exceeds the highest attainable Shannon capacity by any symbol-by-symbol receiver is often termed superadditivity. This phenomenon is also a special case of the well-known quantum mechanical principle called Quantum Nonlocality without Entanglement [5].
Various operational scenarios drive a laser-based communication system into the photon-starved regime, where joint-detection receivers are advantageous. One example is deep-space optical communication (Fig. 1a), where the transmitter laser’s peak power is limited by satellite payload and power, and the aperture dimensions on both ends lead to a significant diffraction-limited loss over the long-range [6, 7].
@@
In this paper, we report the realization of a joint-detection receiver called the optical Green Machine, which is capable of achieving superadditive capacity with the BPSK Hadamard codewords. BPSK encoding is capacity-optimal in the photon-starved regime with bit per photon (bpp) up to two leading terms, approaching the Holevo-limited PIE of an unrestricted encoding [8]. However, the best-known symbol-by-symbol receivers for BPSK decoding can only achieve PIE of bpp. The Green Machine, first proposed as an -port linear-optical interferometer [9], transforms length- BPSK Hadamard codewords into -ary Pulse Position Modulation (PPM) codewords prior to detection (Fig. 1c). Therefore, the achievable capacity of the Green Machine follows the PPM performance in the photon-starved regime bpp [3], which can surpass the best symbol-by-symbol receivers for decoding BPSK (2.89 bpp). Moreover, it reduces the transmitter’s peak power requirement by a factor of compared to PPM.
We build a fiber-based optical Green Machine to realize the first superadditive communication system [9], and thereby demonstrate ‘nonlocality without entanglement’ [5] among a sequence of separated coherent states. The PIE achieved by decoding the length-16 BPSK Hadamard codewords with our Green Machine outperforms ideal homodyne detection receivers after backing out losses. The demonstrated PIE also exceeds the best possible symbol-by-symbol receivers for distinguishing BPSK modulated signals such as the Dolinar receiver, which can obtain the minimum error probability allowed by the laws of quantum mechanics [10, 11, 12, 13, 14, 15, 16].
We further verify that our Green Machine receiver is robust against channel phase noise since the first BPSK symbol can serve as a phase reference. This is a feature unique to the Hadamard encoding paired with the Green Machine, which makes its superadditive capacity even more prominent in noisy-phase environments, e.g., caused by imperfect adaptive optics in a turbulent channel or platform vibrations.
.3 Scalable Optical Green Machine
@@
In the first stage (S1), the four pairs of symbols that are duration apart will interfere. These parallel processes are realized by switching the latter four symbols to the lower path and delaying the first four symbols in the upper path by . The switch is realized by the electro-optical modulator driven by square waves. Then, symbols separated by in time can arrive at the beamsplitter simultaneously for interference. At the end of the first stage (S1), the lower path is delayed by to avoid the degeneracy in temporal modes between the upper and lower paths. The second and third stages have the same structure, but different delays and . Therefore, symbols that are and apart can interfere in the second and third stages, respectively. After the three stages, the input BPSK Hadamard codeword with eight symbols is transformed into an eight-ary PPM codeword, which can be decoded with single-photon detectors by reading the photon arrival time. The complete transformation process of all codewords are depicted in Fig. 1c (more details in Supplementary Information).
.4 Superadditive Communication Demonstration
We then implement our optical Green Machines for coherent optical communications with BPSK Hadamard encoding between the transmitter and receiver (Alice and Bob in Fig. 2a). We test GM1, GM2, GM3, and GM4 for codeword length =2, 4, 8, and 16, respectively. A variable attenuator in the channel is used to control the received mean photon number per symbol. At each channel loss, we repeat multiple experiments to benchmark the performances of the Green Machines with all codewords.
@@
Additionally, the Green Machine receiver does not need any local oscillator (LO), as it utilizes the first symbol of the Hadamard codeword as a self-referenced phase reference to interfere with the following symbols [23, 9]. This feature gives the Green Machine robustness to counter channel-induced phase drift. To validate this benefit, we implement an artificial phase drift in the channel and test the performance of GM3 under phase drift (Fig. 3a). We scan the mean photon numbers with a channel phase drift rate of 100 kilocycles per second (Fig. 3b). Then, we select two and scan the phase drift rate from cycles/s up to million cycles/s. The PIEs of GM3 attained with slow to fast phase drift are exhibited in Fig. 3c and 3d, where the solid curves are fitted by an empirical low-pass band function (Fig. 3c: , MHz, and ; Fig. 3d: , MHz, and ). The performances of the ideal symbol-by-symbol homodyne and Dolinar receivers with no predictive phase tracking are fitted from the Monte Carlo simulations with identical data acquisition time and modulation bandwidth to the Green Machine experiments (total 2.5 million symbols, see Supplementary Information). The heterodyne upper bound assumes perfect I/Q measurements over every symbol.
@@
In summary, we design and build a compact Green Machine receiver with superadditivity demonstrated at low mean photon numbers over a phase-drifting channel. The realized parallel processing of interfering the -symbol Hadamard codewords into -ary PPM codewords through stages, first proposed in Ref. [18, 19], is crucial to the long-term scalability of this temporal Green Machine receiver. Our results represent a significant stepping stone towards attaining a reliable communication rate at the ultimate Holevo bound of channel capacity.
@@
Future directions include lowering crosstalk, adding more stages to the Green Machine, and investigating ways to introduce non-classical optical transforms such as squeezing within the receiver. Eliminating or significantly reducing the losses within the receiver would be particularly important to achieve uncompromised superadditivity, i.e., superadditive capacity without backing out receiver-internal losses. The net loss of the current GM4 system is around dB, primarily introduced by commercial electro-optical modulators (EOMs). With improved coupling efficiency and device quality, the EOM loss could be made negligible [25, 26]. If next-generation EOMs [27] coupled with ultra-low-loss integrated delay lines [28] can reduce the loss to under dB per stage, scaling up to five stages will allow reaching the uncompromised superadditive of bpp, and retain immunity to phase noise in a realistic operational environment (see Supplementary Information). Another direction is to increase the modulation bandwidth towards GHz to compete with the state-of-the-art spectral efficiencies [29], which in turn will reduce the lengths of the delay lines to sub-meter levels. This will help further reduce the size, loss, and design difficulty for a photonic-integrated Green Machine. The higher modulation bandwidth will also help further suppress the noise photons per modulated symbol. Finally, multiplexing single-photon detectors could help overcome the performance limit caused by a single detector’s dead time.
@@
However, the pursuit of quantum receivers that achieve the Holevo capacity is far from being concluded with the Green Machine. Over the decades, many theoretical proposals have elucidated genuine joint-detection receivers capable of achieving superadditivity [30, 31, 32, 33, 9, 34], but practical solutions other than the Green Machine remain elusive. Moreover, there is strong theoretical evidence suggesting that joint-detection receivers describable by the semi-classical theory of photodetection—despite being superadditive-communication capable—cannot achieve the Holevo capacity [35, 12]. Instead, there are proposals of receivers that can achieve the Holevo capacity, e.g., (1) quantum polar code and successive cancellation receiver [3], (2) the slicing receiver [36], (3) codeword unambiguous state discrimination [37], and (4) the sequential-decoding receiver [38] using the multimode collective vacuum or not measurement [39]. The primary bottleneck in their practical realization is the need for high-fidelity or fault-tolerant quantum operations on optically-encoded information [40]. Therefore, in the noisy intermediate-scale quantum (NISQ) era, the Green Machine receiver is undoubtedly a competitive near-term solution for photon-starved communications, which could afford a practical advantage over existing receivers.
@@
Furthermore, the optical Green Machine we have built also has other applications, e.g., in (1) entanglement-assisted classical communications [41], (2) reading information efficiently from a passively-encoded reflective classical memory [42], and (3) loading the single-photon state across time bins into quantum memories [43].
.6 Acknowledgments
The authors acknowledge the support from NASA (80NSSC22K1030) and the NSF ERC Center for Quantum Networks (EEC-1941583). CC and JP thank Allison Rubenok, Khanh Kieu, and Lam Nguyen for their assistance with PM-fiber splicing, thank Rohan Bali for assistance on FPGA, and thank Yiyun Wu and Samar Choura for helping with the thermally stabilized optical breadboard and 3D-printed parts. SG thanks Konrad Banaszek for comments on the performance of homodyne and heterodyne detection receivers with soft information processing.
.7 Author Contributions Statement
All authors conceived the idea and wrote the manuscript together. CC, JP, and LF designed the experiment. CC and JP conducted the experiments and analyzed the data. CC, JP, and SG derived the theoretical model. LF, BS, and SG supervised the work.
.8 Competing Interests Statement
The authors are filing a provisional patent for this receiver design.
.9 Methods
@@
Fiber-optical platform and control. We implement polarization-maintaining fibers and components throughout the system at 1550 nm. The setup is mounted to an optical breadboard with a temperature sensor, an evenly distributed heater, and a PID controller loop to reduce phase noise caused by fiber motion and temperature fluctuation. The delay lines are 16 m for S1, 8 m for S2, 4 m for S3, and 2 m for S4. GM1 is demonstrated using S3; GM2 is composed of S2 followed by S3; GM3 is composed of S1, S2, and S3. For GM1, GM2, and GM3, the symbol time bin duration ns. GM4 is composed of S1, S2, S3, and S4 with ns.
@@
Interspersed active phase correction. The crosstalk is only sensitive to the relative phase drift of two fiber links between the switch and the beamsplitter in each stage [20]. The active phase correction is completed before Alice starts sending the encoded codewords. To correct the phase of each stage, we launch a bright probe laser light (Newfocus Telecom Laser 6428) and tap 1% after each stage’s beamsplitter. We scan the relative phase between symbols from 0 to for times, each lasting a single period of 1 ms, by the PM (Bob) in front of the S1. The scan starts from the first stage. The phase control signal is first generated from an FPGA (RedPitaya STEMlab 125-14), amplified (SRS SIM911 BJT preamplifier), and sent as an amplitude modulation to the square wave signal generator (Rigol DG4102 AWG for S1, S2, and S3, and Rigol DG5102 AWG for S4) connected to the PM. The detector (Newport InGaAs Fiber-Optic Receiver 2011, tuned to 10 kHz bandwidth) reads the power envelope of the varying tapped light, and the output is collected by the same FPGA. The computer fits it by a sine function and triggers the FPGA output voltage to retain maximum or minimum interference. Immediately after, the phase control signal is set to this proper voltage to compensate for the phase difference. Similarly, in series, the phase drifts of all following stages are corrected one by one. The phase correction of each stage takes roughly 10 ms, which is constrained by the Ethernet/Serial port communication latency between the data acquisition units, the control units, and the computer. After phase correction, each relative phase can remain stable for roughly 100 ms in most scenarios.
@@
Testing strategy Alice, channel, and Bob are all controlled by one Python script running on the master computer. The script switches between phase correction and data acquisition. During the phase correction, the channel is switched to zero attenuation by a programmable optical attenuator (JAS Fitel HA9 Optical Attenuator). After phase correction, the channel is switched to a given high attenuation with artificial phase drift to emulate the actual channel. Alice starts encoding a given Hadamard codeword and keeps sending it for 50 ms, while Bob starts recording the photon detection events. If an error occurs during phase correction (e.g. fitting timed out), the data acquisition will be skipped to let the phase correction execute again.

### v4

# Superadditive Communication with the Green Machine as a Practical Demonstration of Nonlocality without Entanglement .1 Abstract

@@
Superadditive Communication with the Green Machine as a Practical Demonstration of Nonlocality without Entanglement
@@
Achieving the ultimate Holevo limit of optical communication capacity requires a joint-detection receiver that makes a collective quantum measurement over multiple modulated symbols. Such superadditivity—a higher communication rate than is achievable by symbol-by-symbol optical detection—is a special case of the well-known nonlocality without entanglement and has yet to be demonstrated. In this article, we propose and demonstrate the design of a joint-detection receiver, the Green Machine, that can achieve superadditivity. We build this receiver, experimentally obtain the transition probability matrix induced by the codebook-receiver pair, and deduce that its capacity surpasses that of any symbol-by-symbol receiver in the photon-starved regime for binary-phase-shift-keying (BPSK)modulation. Our Green Machine receiver can also significantly reduce the transmitter peak power requirement compared with the pulse-position modulation (the conventional modulation format used for deep space laser communication). We further demonstrate that the self-referenced phase makes it resilient to phase noise, e.g., atmospheric turbulence or platform vibrations.
I Introduction
@@
The fundamental limit of reliable optical communication rate, i.e., capacity, is governed by the laws of quantum mechanics and formulated as the Holevo-Schumacher-Westmoreland (HSW) theorem—colloquially known as the Holevo capacity [1, 2]. Holevo capacity is lower bounded by, and often exceeds, the Shannon capacity achievable by any receiver that uses symbol-by-symbol optical detection. This gap becomes more prominent at the photon-starved regime where the received photon flux is low [3, 4]. To close the gap, a joint-detection receiver applying a collective measurement across a codeword comprising multiple symbols is necessary. For a given encoding format, the fact that joint detection can achieve a bits-per-symbol rate that exceeds the highest attainable Shannon capacity by any symbol-by-symbol receiver is often termed superadditivity. This phenomenon is also a special case of the well-known quantum mechanical principle called Quantum Nonlocality without Entanglement [5].
Various operational scenarios drive a laser-based communication system into the photon-starved regime, where joint-detection receivers are advantageous. One example is deep-space optical communication (Fig. 1a and Supplementary Note 2), where the transmitter laser’s peak power is limited by satellite payload and power, and the aperture dimensions on both ends lead to a significant diffraction-limited loss over the extreme distances [6, 7].
@@
In this paper, we report the realization of a joint-detection receiver called the optical Green Machine, which is capable of achieving superadditive capacity paired with the BPSK Hadamard codebook. BPSK modulation is capacity-optimal in the photon-starved regime with bit per photon (bpp) up to two leading terms, approaching the Holevo-limited PIE of an unrestricted encoding [8]. However, the best-known symbol-by-symbol receivers for BPSK decoding can only achieve PIE of bpp. The Green Machine, first proposed as an -port linear-optical interferometer [9], transforms length- BPSK Hadamard codewords into -ary Pulse Position Modulation (PPM) codewords prior to photon detection (Fig. 1c). Therefore, the achievable capacity of the Green Machine follows the PPM performance in the photon-starved regime bpp [3], which can surpass the best symbol-by-symbol receivers for decoding BPSK (2.89 bpp). Moreover, it reduces the transmitter’s peak power requirement by a factor of compared to PPM.
We build a fiber-based optical Green Machine first to realize the superadditive optical communication system [9], and thereby demonstrate nonlocality without entanglement [5] among a sequence of separated weak coherent states. The PIE achieved by decoding all the length-16 BPSK Hadamard codewords with our four-stage Green Machine outperforms ideal homodyne detection receivers after backing out losses. The demonstrated PIE also exceeds the best possible symbol-by-symbol receivers for distinguishing BPSK modulated signals, such as the Dolinar receiver, which can obtain the minimum error probability allowed by the laws of quantum mechanics [10, 11, 12, 13, 14, 15, 16].
We further verify that our Green Machine receiver is robust against channel phase noise by using the leading BPSK symbol as a common pilot phase reference across all codewords. This feature is inherent to the Hadamard encoding paired with the Green Machine, which makes its superadditive capacity even more prominent in noisy-phase environments, such as those caused by imperfect adaptive optics in a turbulent channel or platform vibrations.
II Results
II.1 Scalable Optical Green Machine
@@
In the first stage (S1), the four pairs of symbols that are duration apart will interfere. These parallel processes are realized by switching the latter four symbols to the lower path and delaying the first four symbols in the upper path by . The switch is realized by the electro-optical modulator driven by square waves. Then, symbols separated by in time can arrive at the beamsplitter simultaneously for interference. At the end of the first stage (S1), the lower path is delayed by to avoid the degeneracy in temporal modes between the upper and lower paths. The second and third stages have the same structure, but different delays and . Therefore, symbols that are and apart can interfere in the second and third stages, respectively. After the three stages, the input BPSK Hadamard codeword with eight symbols is transformed into an eight-ary PPM codeword, which can be decoded with single-photon detectors by reading the photon arrival time. The complete transformation process of all codewords is depicted in Fig. 1c.
II.2 Superadditive Communication Demonstration
We then implement our optical Green Machines for coherent optical communication with BPSK Hadamard encoding between the transmitter and receiver (Alice and Bob in Fig. 2a). We test GM1, GM2, GM3, and GM4 for codeword length =2, 4, 8, and 16, respectively. A variable attenuator in the channel is used to control the received mean photon number per symbol. At each channel loss, we repeat multiple experiments to benchmark the performances of the Green Machines with all codewords.
@@ II.2 Superadditive Communication Demonstration
Additionally, the Green Machine receiver does not need any local oscillator (LO), as it utilizes the first symbol of the Hadamard codeword as a self-referenced phase reference to interfere with the following symbols [23, 9]. This feature gives the Green Machine robustness to counter channel-induced phase drift. To validate this benefit, we implement an artificial phase drift in the channel and test the performance of GM3 under phase drift (Fig. 3a). We scan the mean photon numbers with a channel phase drift rate of 100 kilocycles per second (Fig. 3b). Then, we select two and scan the phase drift rate from cycles/s up to million cycles/s. The PIEs of GM3 attained with slow to fast phase drift are exhibited in Fig. 3c and 3d, where the solid curves are fitted by an empirical low-pass band function (Fig. 3c: , MHz, and ; Fig. 3d: , MHz, and ). The performances of the ideal symbol-by-symbol homodyne and Dolinar receivers with no predictive phase tracking are fitted from the Monte Carlo simulations with identical data acquisition time and modulation bandwidth to the Green Machine experiments (total 2.5 million symbols, highlighted in Supplementary Note 5A). The heterodyne upper bound assumes perfect I/Q measurements over every symbol.
@@ III Discussion
In summary, we design and build a compact optical Green Machine receiver with superadditivity demonstrated at low mean photon numbers over a phase-drifting channel. The realized parallel processing of interfering the -symbol Hadamard codewords into -ary PPM codewords through stages, first proposed in Ref. [18, 19], is crucial to the long-term scalability of this temporal Green Machine receiver. Our results represent a significant stepping stone towards attaining a reliable communication rate at the ultimate Holevo bound of channel capacity. So far, the communication system we have implemented only contains the optical realization of a pulsed laser transmitter, a fiber channel with loss and phase noise, and a receiver capable of realizing superadditive communication capacity. To demonstrate a fully functional communication system that can send messages reliably at a superadditive rate, one would need to incorporate an appropriate forward error correction code to correct the errors introduced by the superchannel induced by our Hadamard code Green Machine receiver combination—quantified by our measured transition probability matrix—such as the Reed-Solomon code or a low-density parity-check (LDPC) code, and an FPGA system that realizes the encoding and decoding operations.
@@
Future directions include lowering crosstalk, adding more stages to the Green Machine, and investigating ways to introduce non-classical optical transforms, such as squeezing within the receiver. Eliminating or significantly reducing the losses within the receiver would be particularly important to achieve uncompromised superadditivity, i.e., superadditive capacity without backing out receiver-internal losses. The net loss of the current GM4 system is around dB, primarily introduced by commercial electro-optical modulators (EOMs). With improved coupling efficiency and device quality, the EOM loss could be made negligible [25, 26]. If next-generation EOMs [27] coupled with ultra-low-loss integrated delay lines [28] can reduce the loss to under dB per stage, scaling up to five stages will allow reaching the uncompromised superadditive of bpp, and retain immunity to phase noise in a realistic operational environment (see Supplementary Note 3). Another direction is to increase the modulation bandwidth towards GHz to compete with the state-of-the-art spectral efficiencies [29], which in turn will reduce the lengths of the delay lines to sub-meter levels. This will help further reduce the size, loss, and design difficulty for a photonic-integrated Green Machine. The higher modulation bandwidth will also help further suppress the noise photons per modulated symbol. Finally, multiplexing single-photon detectors could help overcome the performance limit caused by a single detector’s dead time.
@@
However, the pursuit of quantum receivers that achieve the Holevo capacity is far from being concluded with the Green Machine. Over the decades, many theoretical proposals have elucidated genuine joint-detection receivers capable of achieving superadditivity [30, 31, 32, 33, 9, 34], but practical solutions other than the Green Machine remain elusive. Moreover, there is strong theoretical evidence suggesting that joint-detection receivers describable by the semi-classical theory of photodetection—despite being superadditive-communication capable—cannot achieve the Holevo capacity [35, 12]. Instead, there are proposals of receivers that can achieve the Holevo capacity, e.g., (1) quantum polar code and successive cancellation receiver [3], (2) the slicing receiver [36], (3) codeword unambiguous state discrimination [37], and (4) the sequential-decoding receiver [38] using the multimode collective Vacuum or Not Measurement [39]. The primary bottleneck in their practical realization is the need for high-fidelity or fault-tolerant quantum operations on optically-encoded information [40]. Therefore, in the noisy intermediate-scale quantum (NISQ) era, the Green Machine receiver is undoubtedly a competitive near-term solution for photon-starved communication, which could afford a practical advantage over existing receivers.
@@
Furthermore, the optical Green Machine we have built also has other applications, e.g., in (1) entanglement-assisted classical communication [41], (2) reading information efficiently from a passively-encoded reflective classical memory [42], and (3) loading the single-photon state across time bins into quantum memories [43].
IV Methods
IV.1 Fiber-optical platform and control.
@@ IV Methods
We implement polarization-maintaining fibers and components throughout the system at 1550 nm. The setup is mounted to an optical breadboard with a temperature sensor, an evenly distributed heater, and a PID controller loop to reduce phase noise caused by fiber motion and temperature fluctuation. The delay lines are 16 m for S1, 8 m for S2, 4 m for S3, and 2 m for S4. GM1 is demonstrated using S3; GM2 is composed of S2 followed by S3; GM3 is composed of S1, S2, and S3. For GM1, GM2, and GM3, the symbol time bin duration ns. GM4 is composed of S1, S2, S3, and S4 with ns.
@@ IV.2 Interspersed active phase correction
The crosstalk is only sensitive to the relative phase drift of two fiber links between the switch and the beamsplitter in each stage [20]. The active phase correction is completed before Alice starts sending the encoded codewords. To correct the phase of each stage, we launch a bright probe laser light (Newfocus Telecom Laser 6428) and tap 1% after each stage’s beamsplitter. We scan the relative phase between symbols from 0 to for times, each lasting a single period of 1 ms, by the PM (Bob) in front of the S1. The scan starts from the first stage. The phase control signal is first generated from an FPGA (RedPitaya STEMlab 125-14), amplified (SRS SIM911 BJT preamplifier), and sent as an amplitude modulation to the square wave signal generator (Rigol DG4102 AWG for S1, S2, and S3, and Rigol DG5102 AWG for S4) connected to the PM. The detector (Newport InGaAs Fiber-Optic Receiver 2011, tuned to 10 kHz bandwidth) reads the power envelope of the varying tapped light, and the output is collected by the same FPGA. The computer fits it by a sine function and triggers the FPGA output voltage to retain maximum or minimum interference. Immediately after, the phase control signal is set to this proper voltage to compensate for the phase difference. Similarly, in series, the phase drifts of all following stages are corrected one by one. The phase correction of each stage takes roughly 10 ms, which is constrained by the Ethernet/Serial port communication latency between the data acquisition units, the control units, and the computer. After phase correction, each relative phase can remain stable for roughly 100 ms in most scenarios.
IV.3 Testing strategy
@@
Alice, channel, and Bob are all controlled by one Python script running on the master computer. The script switches between phase correction and data acquisition. During the phase correction, the channel is switched to zero attenuation by a programmable optical attenuator (JAS Fitel HA9 Optical Attenuator). After phase correction, the channel is switched to a given high attenuation with artificial phase drift to emulate the actual channel. Alice starts encoding a given Hadamard codeword and keeps sending it for 50 ms, while Bob starts recording the photon detection events. If an error occurs during phase correction (e.g., fitting timed out), the data acquisition will be skipped to let the phase correction execute again.
IV.4 Data acquisition